OpenStudy (mr.math):

Find the sum \(\large \sum_{k=1}^{n} k!(k^2+k+1).\)

OpenStudy (anonymous):

\( \large \sum \limits_{k=1}^{n} k!(k^2+k+1) = (n+1)!(n+1)-1 \) PS: Latex Tip included ;-)

OpenStudy (mr.math):

You should show work, so others can learn :-)

OpenStudy (zarkon):

\[k^2+k+1=(k+1)^2-k\]

OpenStudy (zarkon):

is my hint good enough?

OpenStudy (mr.math):

Yes.

OpenStudy (anonymous):

One more hint, \( \large \sum \limits_{k=1}^{n} k\times k! = (n+1)! -1\)

OpenStudy (anonymous):

oops MR.Math is great :)

OpenStudy (mr.math):

lol, I already know the answer. I just wanted to post something, and I think it's a nice problem.

OpenStudy (anonymous):

haha :)

OpenStudy (anonymous):

Mr.Math here: http://openstudy.com/study#/updates/4f2ebf12e4b0571e9cbaf294 ;-)